OpenStudy (anonymous):

6(5 + 3) = 6 • 5 + 6 • y what is the value of y?

OpenStudy (anonymous):

Here let's do the info we can solve right now :) 6(5+3) = 6*5+6*y 6(8) = 30 + 6*y Can you tell me what is 6*8?

OpenStudy (kaydra14):

6*8 is 48

OpenStudy (anonymous):

Yes, but please let the asker answer :P So now we know that 48 = 30 + 6*y Now we must subtract 30 on both of the equation:) 48 - 30 = 30 - 30 + 6*y 48 - 30 = 6*y Can you tell me what is 48-30?

OpenStudy (kaydra14):

so you dont want me to answer it @sammixbib

OpenStudy (anonymous):

48-30= 18

OpenStudy (anonymous):

I want you to help MathNerdie with the question, but don't answer what I am asking him/her to solve :) I am trying to show him/her step by step how to do this problem and he/she would get the best help if he/she answered it

OpenStudy (anonymous):

so it would be 6x18?

OpenStudy (anonymous):

Corect! Now we know 18 = 6*y now the last thing we need to do is divide 6 on both sides of the equation 18/6 = 6*y/6 18/6 = y Can you tell me what is 18/6? Once you figure that out, that is your answer!

OpenStudy (anonymous):

No, 18/6 :)

OpenStudy (anonymous):

Aka 18 divided by 6

OpenStudy (anonymous):

it would be 3?

TheSmartOne (thesmartone):

Yes it is 3... lol @sammixbib You left them... lol

OpenStudy (anonymous):

y=3 -6y+18=0